OpenStudy (sloppycanada):

In January of 2010, the population of a small town called Riley was 1,250. Its population has been decreasing at a steady rate of 8% per year. According to this model, what was Riley's population on January 2000?

OpenStudy (sloppycanada):

I know right away that I use the following formula - A = P * e^(r)(t)) this is the formula for continuous growth or decay. In our case it's decay.

OpenStudy (sloppycanada):

From there, I'm not quite sure what I want to do. Do I set up the equation as the following - A = 1250 * e^(-.08)(-10)

OpenStudy (sloppycanada):

Am I correct on my thinking or...?

jimthompson5910 (jim_thompson5910):

``` I know right away that I use the following formula - A = P * e^(r)(t)) this is the formula for continuous growth or decay ``` sorry that's not correct. The population isn't decreasing at a continuous rate. It's only decreasing on an annual basis

jimthompson5910 (jim_thompson5910):

You actually use this formula \[\Large A = P(1+r)^t\] where A = final amount after t years P = initial amount r = growth or decay rate if r > 0 then we have growth, if r < 0 then we have decay t = time in years

OpenStudy (mathstudent55):

I would have thought that you use a yearly compounding formula. Also, the initial value in 2010 is unknown. The future value is 1250 in 2010.

jimthompson5910 (jim_thompson5910):

`In January of 2010, the population of a small town called Riley was 1,250. Its population has been decreasing at a steady rate of 8% per year. According to this model, what was Riley's population on January 2000? ` in this case, A = 1250 P = unknown r = 0.08 t = 10 use these values and the formula to solve for P

jimthompson5910 (jim_thompson5910):

sorry I meant to make r negative. It should be r = -0.08

OpenStudy (sloppycanada):

But as a hint, it says to use the continuously compounding formula?

jimthompson5910 (jim_thompson5910):

You're right, the key phrase `at a steady rate of ` is why you use the continuously compounding formula. My bad

jimthompson5910 (jim_thompson5910):

somehow I missed that too

OpenStudy (sloppycanada):

Wouldn't t = 10? Since it's the time between 2000 and 2010?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (sloppycanada):

Alright, so using the continuous growth or decay formula, I'm looking at 1250/(e^(-.08)(10) = P?

OpenStudy (sloppycanada):

But if I solve things like that, I strangely get a number greater than 1250. Which makes me think that I'm using the wrong values for P and A.

OpenStudy (mathstudent55):

If the population decreased steadily for 10 years and ended up at 1250 in 2010, then it must have been greater than 1250 to start with in the year 2000. You are looking for a number greater than 1250.

OpenStudy (sloppycanada):

Okay, good. Right, forgot about that. So in 2000, the population was about 2782

OpenStudy (mathstudent55):

Sounds reasonable.

jimthompson5910 (jim_thompson5910):

As you can see, the graph slowly goes down hill following along a curve of sorts. See attached So yes, as said earlier, the initial amount should be larger

OpenStudy (mathstudent55):

\(\large A = Pe^{rt} \) \(\large A = 2782 e^{-0.08 \times 10} \) \(\large A = 1250.033178\)

OpenStudy (nincompoop):

You should get a larger population in 2010 than in 2010 if the rate decreases every year even for whatever model it is. It should automatically prompt you that your slope is going to be negative so that means the population is more in the past than in the future in this case.

OpenStudy (sloppycanada):

Thanks for the help!

OpenStudy (mathstudent55):

@nincompoop You wrote above: "You should get a larger population in 2010 than in 2010" ???

OpenStudy (nincompoop):

2000 than in 2010

OpenStudy (nincompoop):

my badness typoness